# Repeatable Energy-Efficient Perching for Flapping-Wing Robots Using Soft Grippers

### v1

@@ Abstract
With the emergence of new flapping-wing micro aerial vehicle (FWMAV) designs, a need for extensive and advanced mission capabilities arises. FWMAVs try to adapt and emulate the flight features of birds and flying insects. While current designs already achieve high manoeuvrability, they still almost entirely lack perching and take-off abilities. These capabilities could, for instance, enable long-term monitoring and surveillance missions, and operations in cluttered environments or in proximity to humans and animals. We present the development and testing of a framework that enables repeatable perching and take-off for small to medium-sized FWMAVs, utilising soft, non-damaging grippers. Thanks to its novel active-passive actuation system, an energy-conserving state can be achieved and indefinitely maintained while the vehicle is perched. A prototype of the proposed system weighing under 39 g was manufactured and extensively tested on a 110 g flapping-wing robot. Successful free-flight tests demonstrated the full mission cycle of landing, perching and subsequent take-off. The telemetry data recorded during the flights yields extensive insight into the system’s behaviour and is a valuable step towards full automation and optimisation of the entire take-off and landing cycle.
Keywords Grippers and Other End-Effectors Actuation and Joint Mechanisms Biologically-Inspired Robots Soft Robotics Aerial Systems: Applications Mechanism Design
@@ 1 Introduction
Bioinspired flapping-wing micro aerial vehicles (FWMAVs) fill in the gap between fixed-wing and multirotor MAVs. They are highly agile and efficient at low flight speeds and small scales, and their extensive flight envelope encompasses hover, glide and cruise flight modes. On top of this, FWMAVs are inherently safe, thanks to their lightweight structures and a propulsion system based on lightweight, often flexible wings rather than hazardous propellers. These features make FWMAVs highly promising for applications in cluttered spaces and delicate environments, especially near humans, flora or fauna. Even so, these robots are barely used outside research settings due to their high complexity and limited flight time and payload. Most FWMAVs can only fly for a few minutes (Wu et al.,, 2023) while carrying little or no useful payload and need to be launched and retrieved manually (Hammad and Armanini,, 2024). All this makes complex missions unthinkable without significant improvements.
One approach to enable longer-duration and more complex missions that has been researched extensively in the multirotor community is perching (e.g. (Culler et al.,, 2012; Doyle et al.,, 2013; Popek et al.,, 2018; Sato et al.,, 2021)), i.e. being able to attach to existing structures in the surroundings. Culler et al., (2012) and Doyle et al., (2013), for instance, proposed passive, compliant gripping mechanisms for perching on circular or square tubes and demonstrated successful perching of quadrotors in manual guided tests; however, free-flight testing was lacking. Popek et al., (2018) demonstrated an active-passive grasping mechanism for multirotor MAVs, achieving, for the first time, automatic free-flight perching capabilities on circular bars, relying solely on onboard sensing and guidance. Sato et al., (2021) developed a passive gripping mechanism for multirotor UAVs, enabling them to perch on the upper surface of vertical cylindrical objects, which was shown in manually steered flight tests. Over the past few years, numerous studies have investigated different types of perching solutions for multirotors, as reviewed, for instance, in Roderick et al., (2017) and Meng et al., (2022).
@@
Only recently, a small number of studies have considered the challenge of perching with flapping-wing robots (Broers and Armanini,, 2022; Gomez-Tamm et al.,, 2020; Graule et al.,, 2016; Zufferey et al.,, 2022). Graule et al., (2016) achieved perching and take-off utilising an active electro-adhesive pad on a tethered small-insect scale flapping-wing robot. The robot, with its attachment pads, weighed 100 mg and was able to attach to flat undersides of different materials. Gomez-Tamm et al., (2020) developed a shape memory alloy (SMA) activated grasping mechanism for integration onto larger FWMAVs. They demonstrated gripping and static perching capabilities, as well as take-off from a perched position; however, free-flight perching was not achieved. Zufferey et al., (2022) recently proposed the first successful active-passive perching system for a larger, tailed FWMAV weighing 700 g. It utilises an actuated leg with an attached bistable claw mechanism that passively closes upon contact with a circular branch; resetting the mechanism is realised via a DC motor and completed within 20 seconds. However, a take-off sequence hasn’t been implemented. (Zufferey et al.,, 2022)
@@
Perching, therefore, has been achieved under certain constraints in the ultra-small and large size class, leaving a gap for the highly interesting and widely studied category of small- to medium-sized FWMAVs. Additionally and importantly for real-life applications, the full free-flight perching manoeuvre cycle of at-will repeatable landing, perching and take-off has seldom been demonstrated. The flapping-wing robot by Graule et al., (2016), for instance, represents a successful candidate. However, their approach only physically works for ultralight FWMAVs.
@@
Even in the more extensively studied field of multirotor and fixed-wing aerial robots, only a few were able to demonstrate the required full capabilities. From the previously mentioned examples, Popek et al., (2018) only demonstrated the landing, while Sato et al., (2021) achieved repeated landing and take-off, however, utilising a simpler, solely passive attachment design.
@@
In this paper, we present a simple, energy-efficient and non-damaging perching framework for FWMAVs, allowing repeatable perching and take-off cycles during a mission. Building on a soft gripping mechanism developed in our previous research (Broers and Armanini,, 2022) and tested on a multirotor platform, we have devised and implemented a system that enables perching functionality on a representative medium-sized (100 g - 150 g) tailless FWMAV platform (see Fig. 1).
@@
A central element of our framework is a new, adapted and improved design of the previously introduced soft gripper concept. Key advancements with respect to the original concept include an overhaul of the mechanical design, leading to significant weight reduction as well as increased capabilities and robustness. Finally, through an integration of the perching mechanism onto a flapping-wing system, extensive tests were conducted, demonstrating our approach’s performance and viability, as well as establishing its operation envelope.
@@
Demonstration of the complete sequence of landing, perching, and take-off in a free-flight environment; and validation of the robustness and reliability of the concept through repeated consecutive flight tests. The successful display of a take-off from a perched position, in particular, is frequently neglected in other studies (e.g. (Popek et al.,, 2018; Zufferey et al.,, 2022)).
Analysis of data from comprehensive flight and gripping tests of an integrated system to help improve gripper-based perching systems and provide data on a realistic performance envelope.
@@ 2 Design and Integration
Our perching framework is based on a design concept for a gripping mechanism established in Broers and Armanini, (2022). The main focus of the aforementioned study was the mechanism itself rather than its application for FWMAV perching. While ultimately intended for flapping-wing vehicles, the mechanism was previously only tested on a multirotor and not yet fully adapted for the FWMAV case. In this work, the same basic design concept was further developed into a complete perching framework for use in an exemplary FWMAV platform. This included specifically adapting the mechanism design for usage on FWMAVs, fully integrating the mechanism on a FWMAV test platform, and extensively testing all operation and flight phases of the resulting perching framework. The resulting design and its components can be seen in Fig. 2.
To demonstrate the functionality of our system, we used a commercially available robot as a test platform, i.e. the “Flapper Nimble+”, produced by Flapper Drones, (2024) and based on the DelFly Nimble developed by Karásek et al., (2018).
The gripping mechanism used in our perching framework was developed using a systematic biomimetic design process proposed by Fayemi et al., (2017) and relies on three main components:
@@
A detailed discussion of the gripping mechanism our perching framework is based on can be found in Broers and Armanini, (2022). Only a concise overview is provided here for completeness, while all the novel elements of our current design are discussed in detail.
@@ 2.1 Mechanism Design
The gripping mechanism, with its components depicted in Fig. 2, is comprised of an actuated linkage that drives the attached grippers to enclose the desired perching object. This bioinspired linkage mechanism consists of four bars connected through revolute joints, with the top bar fixed to the flying platform. A gripper is attached to one of the outer bars and the middle bar. The characteristic part of this four-bar linkage design is that it decouples closing and opening forces, which can be implemented by pulling on one of the outer bars that are connected through the middle bar. For example, applying a closing force results in a scissor-like motion of the linkage, pushing the attached grippers together. On the other hand, applying an opening force straightens the opposing outer and middle bars, making them parallel.
To integrate the gripping mechanism on a generic tailless FWMAV, the top bar is used as a central connection point to the underside of the flying platform. To accommodate various approach paths, a folding symmetry was achieved by introducing a kink in the top bar, mirroring the outer bar, which enables the opening functionality. The bars were 3D-printed from Tough PLA to be lightweight and to provide strong resistance to bending forces, while the linking joints were each made from a polymer sleeve bearing in combination with a hollow brass axle for good friction characteristics.
The soft grippers work according to the Fin Ray effect, which can be observed in fish fins and was first found by Leif Kniese (Pfaff et al.,, 2011). The triangular structure of the fin-inspired grippers, containing cross-struts, reacts to a contact force in a compliant manner, bending towards the direction of the impacting force. The basic layout of the grippers was adapted from Broers and Armanini, (2022) and 3D printed from TPU. The geometry consists of a flexible outer shape representing an extruded tall isosceles triangle containing rigid cross-struts hinged through flexible connections to the outer hull. This allows for the desired compliant bending behaviour and soft design, preventing damage to both the perching object and the gripping mechanism while also maximising contact area for various perching object sizes and accommodating irregular surfaces. An attachment to the linkage through a slot system was devised for easy assembly and change of grippers. This leads to increased gripper stability while reducing weight. To optimise the attachment properties of the grippers to the perching object, their inner surface was lined with a thin sandpaper-like anti-slip tape.
@@ 2.2 Actuation Design
The gripping mechanism is operated in an active-passive manner, similar to how many birds perch. This approach allows for a high achievable gripping power through an active perching mechanism while conserving energy in a perched state through passive engagement. In birds’ feet, this is facilitated by the digital tendon locking mechanism, which acts like a ratchet, where the tendon interlocks with a tendon sheath to maintain a perched state (Quinn and Baumel,, 1990).
A servo motor with a power output of 55 Ncm torque at 7.4 V was chosen for the active actuation.
For an optimal centre of gravity location, i.e. centred on the Flapper’s vertical axis, the servo-motor was integrated into the middle of the top bar.
A lever with three arms was designed for the servo-motor, with a pull cable attached to each of the three levers (see Fig. 2, made from sturdy and lightweight nylon fishing rod:
The right lever connects to one of the outer bars to enable the closing of the gripping mechanism once the servo-motor turns counterclockwise.
@@
The left lever connects to the other outer linkage bar and facilitates the opening of the gripping mechanism once the servo-motor turns clockwise.
For the passive part, i.e. allowing sustained perching without energy expenditure, a ratchet mechanism had been proposed in our previous work (Broers and Armanini,, 2022). However, its rudimentary design could only achieve low precision and, crucially, did not allow for a resetting functionality, thus making an automated take-off from a perched condition impossible.
In the new design, significant changes were introduced to address these limitations, and a new custom spring-loaded ratchet mechanism was developed. The new mechanism consists of a ratchet wheel connected to one of the moving linkage bars and an enclosed spring-loaded pawl that is affixed to the top bar. The ratchet wheel was precision CNC-machined from aluminium, enabling good wear resistance and tight tolerances. Its geometry was improved with miniaturised gear teeth; thus maximising gear teeth count for more precise interlocking. Furthermore, the gear tooth flank design was enhanced for the wheel and pawl, enabling a frictionless tightening of the ratchet. The resulting geometry on the ratchet wheel consisted of 29 evenly spaced teeth over a sector of 37.9° and a root diameter of 64 mm. The tooth profile is an approximate right triangle, with an angle between the perpendicular flank and the hypotenuse of 18.3°. Additionally, a fillet radius of 0.15 mm was chosen for the tooth tip and root for manufacturing purposes.
The curved pawl was likewise CNC-machined from aluminium and included a protruded cable attachment point to enable resetting. It is enclosed in a housing manufactured by precise digital light processing and made out of IGUS Iglidur i3000-PR, a material optimised for sliding parts. Within the housing, a compressed spring constantly pushes the sliding pawl onto the ratchet wheel and therefore allows for the servo-actuated resetting functionality: if the pawl is pulled into the housing, it further compresses the spring and disengages from the ratchet wheel (see Fig. 3).
@@
Actuating the gripping mechanism’s closure and thus grasping a perching object is facilitated through a custom trigger system: A trigger fork establishes contact with the perching object and activates a mechanical switch housed in the top bar of the linkage. This basic design was kept from the previous prototype (Broers and Armanini,, 2022) but improved upon to account for a flapping-wing platform: Its housing features tighter tolerances and a better attachment to the mechanical switch, allowing for more reliable triggering and higher stability in the perched state. Additionally, the required actuation force of the switch was changed from 2.6 N to 1.6 N to accommodate a lighter drone and enable triggering even at lower impact velocities.

### v2

# September 2025 Repeatable Energy-Efficient Perching for Flapping-Wing Robots Using Soft-Grippers

@@ Abstract
With the emergence of new flapping-wing micro aerial vehicle (FWMAV) designs, a need for extensive and advanced mission capabilities arises. FWMAVs try to adapt and emulate the flight features of birds and flying insects. While current designs already achieve high manoeuvrability, they still almost entirely lack perching and take-off capabilities. These capabilities would enable long-term monitoring and surveillance operations, and more complex and multifaceted missions in cluttered environments. We present the development and testing of a framework that enables repeatable perching and take-off for small to medium-sized FWMAVs, utilising soft, non-damaging grippers. Thanks to its novel active-passive actuation system, an energy-conserving state can be achieved and indefinitely maintained while the vehicle is perched. This actuation system is inspired by the digital tendon locking mechanism observed in perching birds and allows for high gripping power and minimal energy usage with a low weight penalty. A prototype of the proposed system weighing under 39 g was manufactured and extensively tested on a 110 g flapping-wing robot. Successful free-flight tests demonstrated the full mission cycle of landing, perching and subsequent take-off. The telemetry data recorded during the flights yields extensive insight into the system’s behaviour and is a valuable step towards full automation and optimisation of the entire take-off and landing cycle.
Keywords: Flapping-Wing MAV (FWMAV), Aerial Robotics, Soft Robotics, Grippers, Perching, Flapping-Wing Take-Off and Landing, Micro Aerial Vehicles (MAVs)
A Preprint
@@ 1 Introduction
Bioinspired flapping-wing micro aerial vehicles (FWMAVs) fill in the gap between fixed-wing and multirotor MAVs. They are highly agile and efficient at low flight speeds and small scales, and their extensive flight envelope encompasses hover, glide and cruise flight modes. On top of this, FWMAVs are inherently safe, thanks to their lightweight structures and a propulsion system based on lightweight, often flexible wings rather than hazardous propellers. These features make FWMAVs highly promising for applications in cluttered spaces and delicate environments, especially near humans, flora or fauna. Even so, these robots are barely used outside research settings due to their high complexity and limited flight time and payload. Most FWMAVs can only fly for a few minutes (Wu et al., 2023) while carrying little or no useful payload and need to be launched and retrieved manually (Hammad and Armanini, 2024). All this makes complex missions unthinkable without significant improvements.
One approach to enable longer-duration and more complex missions that has been researched extensively in the multirotor community is perching (e.g. (Culler et al., 2012; Doyle et al., 2013; Popek et al., 2018; Sato et al., 2021)), i.e. being able to attach to existing structures in the surroundings. Culler et al. (2012) and Doyle et al. (2013), for instance, proposed passive, compliant gripping mechanisms for perching on circular or square tubes and demonstrated successful perching of quadrotors in manual guided tests; however, free-flight testing was lacking. The prototype by Culler et al. (2012) weighed 98 g, while the two-legged mechanism by Doyle et al. (2013) weighed 478 g. Popek et al. (2018) demonstrated an active-passive grasping mechanism for multirotor MAVs, achieving, for the first time, automatic free-flight perching capabilities on circular bars, relying solely on onboard sensing and guidance. Their manipulator weighed 372 g, leading to a system weight of 2.3 kg (Popek et al., 2018). Sato et al. (2021) developed a passive gripping mechanism for multirotor UAVs, weighing 640 g, enabling them to perch on the upper surface of vertical cylindrical objects, which was shown in manually steered flight tests. Over the past few years, numerous studies have investigated different types of perching solutions for multirotors, as reviewed, for instance, in Roderick et al. (2017) and Meng et al. (2022).
@@
Only recently, a small number of studies have considered the challenge of perching with flapping-wing robots (Broers and Armanini, 2022; Gomez-Tamm et al., 2020; Graule et al., 2016; Zufferey et al., 2022). Graule et al. (2016) achieved perching and take-off utilising an active electro-adhesive pad on a tethered small-insect scale flapping-wing robot. The robot, with its attachment pads, weighed 100 mg and was able to attach to flat undersides of different materials. Gomez-Tamm et al. (2020) developed a shape memory alloy (SMA) activated grasping mechanism for integration onto larger FWMAVs. They demonstrated gripping and static perching capabilities, as well as take-off from a perched position; however, free-flight perching was not achieved. Zufferey et al. (2022) recently proposed the first successful active-passive perching system, weighing 184 g for a larger, tailed FWMAV weighing 700 g in total. It utilises an actuated leg with an attached bistable claw mechanism that passively closes upon contact with a circular branch; resetting the mechanism is realised via a DC motor and completed within 20 seconds. However, a take-off sequence hasn’t been implemented. (Zufferey et al., 2022)
@@
Perching, therefore, has been achieved under certain constraints in the ultra-small and large size classes, leaving a gap for the highly interesting and widely studied category of small- to medium-sized FWMAVs. This area is particularly challenging as it cannot rely on the different physical conditions that allow for, e.g. adhesive perching at ultra-light sizes, while still being severely weight-constrained to not be able to integrate powerful actuators or actively maintain its perch due to battery limitations. Additionally and importantly for real-life applications, the full free-flight perching manoeuvre cycle of at-will repeatable landing, perching and take-off has seldom been demonstrated. The ultralight flapping-wing robot by Graule et al. (2016), for instance, represents a successful candidate.
@@
Even in the more extensively studied field of multirotor and fixed-wing aerial robots, only a few were able to demonstrate the required full capabilities. From the previously mentioned examples, Popek et al. (2018) only demonstrated the landing, while Sato et al. (2021) achieved repeated landing and take-off, however, utilising a simpler, solely passive attachment design.
@@
In this paper, we present a simple, energy-efficient and non-damaging perching framework for FWMAVs, allowing repeatable perching and take-off cycles during a mission. Building on a soft gripping linkage mechanism developed in our previous research (Broers and Armanini, 2022) and tested on a multirotor platform, we have devised and implemented a system that enables perching functionality on a representative medium-sized (100 g - 150 g) tailless FWMAV platform (see figure 1).
@@
A central element of our framework is a new, adapted and improved design of the previously introduced soft gripper concept. Key advancements with respect to the original concept include an overhaul of the mechanical design and actuation system, leading to additional capabilities, achieving a full and repeatable perch cycle, as well as weight reductions and performance improvements. Furthermore, through an integration of the perching mechanism onto a flapping-wing system, extensive tests were conducted, demonstrating our approach’s performance and viability, as well as establishing its operation envelope.
@@
Demonstration of the complete sequence of landing, perching, and take-off in a free-flight environment; and validation of the robustness and reliability of the concept through repeated consecutive flight tests. The successful display of a take-off from a perched position, in particular, is frequently neglected in other studies (e.g. in Popek et al. (2018)).
Analysis of data from comprehensive flight and gripping tests of an integrated system to help improve gripper-based perching systems and determine a realistic performance envelope.
@@ 2 Design and Integration
Our perching framework is based on a design concept for a bioinspired gripping mechanism established in Broers and Armanini (2022). The aforementioned study introduced an initial version of the mechanism, enabling perched landing of an MAV. However, the mechanism did not allow for take-off. Moreover, while it was ultimately intended for flapping-wing vehicles, the mechanism was only tested on a multirotor and not yet fully adapted for the FWMAV case. In this work, we build on the same basic design concept and develop it into a complete perching framework for use on an exemplary FWMAV platform. This included specifically adapting the mechanism design for usage on FWMAVs and fully integrating the mechanism on a FWMAV test platform. Furthermore, the perching capabilities were extended to include take-off and to enable repeatability. Extensive testing was performed to demonstrate and characterise all operation and flight phases of the resulting perching framework. The resulting design and its components can be seen in figure 2 with a photo of the produced prototype in figure 3.
To demonstrate the functionality of our system, we used a commercially available robot as a test platform, i.e. the “Flapper Nimble+”, produced by Flapper Drones (2024) and based on the DelFly Nimble developed by Karásek et al. (2018).
The gripping mechanism used in our perching framework was developed using a systematic biomimetic design process proposed by Fayemi et al. (2017) and relies on three main components:
@@
A detailed discussion of the gripping mechanism our perching framework is based on can be found in Broers and Armanini (2022). Only a concise overview is provided here for completeness, while all the novel elements of our current design are discussed in detail.
@@ 2.1 Mechanism Design
The gripping mechanism, with its components depicted in figure 2, is comprised of an actuated linkage that drives the attached grippers to enclose the desired perching object. This bioinspired linkage mechanism consists of four bars connected through revolute joints (as seen in the numbered linkage components in figure 2), with the top bar (4) fixed to the flying platform. The grippers are attached to one of the outer bars (3) and the middle bar (2). The characteristic part of this four-bar linkage design is that it decouples closing and opening forces. For example, applying a closing force at bar (3) results in a scissor-like motion of the linkage, pushing the attached grippers together. On the other hand, applying an opening force at bar (1) straightens the opposing outer (3) and middle (2) bars, making them parallel.
To integrate the gripping mechanism on a generic tailless FWMAV, the top bar (4) is used as a central connection point to the underside of the flying platform. To accommodate various approach paths, a folding symmetry was achieved by introducing a kink in the top bar, mirroring the outer bar (1), which enables the opening functionality. The bars were 3D-printed from Tough PLA to be lightweight and to provide strong resistance to bending forces, while the linking joints were each made from a polymer sleeve bearing in combination with a hollow brass axle for good friction characteristics.
The soft grippers work according to the Fin Ray effect, which can be observed in fish fins and was first found by Leif Kniese (Pfaff et al., 2011). The triangular structure of the fin-inspired grippers, containing cross-struts, reacts to a contact force in a compliant manner, bending towards the direction of the impacting force. The basic layout of the grippers was adapted from Broers and Armanini (2022) and 3D printed from TPU. The geometry consists of a flexible outer shape representing an extruded tall isosceles triangle containing rigid cross-struts hinged through flexible connections to the outer hull. This allows for the desired compliant bending behaviour and soft design, preventing damage to both the perching object and the gripping mechanism while also maximising contact area for various perching object sizes and accommodating irregular surfaces. An attachment to the linkage through a slot system was devised for easy assembly and change of grippers. This leads to increased gripper stability while reducing weight. To optimise the attachment properties of the grippers to the perching object, their inner surface was lined with a thin sandpaper-like anti-slip tape. The gripper’s geometry dictates the range of compatible perching objects. In our study, an opening angle of 35° was chosen, leading to a gripper tip distance of 66 mm.
@@ 2.2 Actuation Design
The gripping mechanism is operated in an active-passive manner, similar to how many birds perch. This approach allows for a high achievable gripping power through an active perching mechanism while conserving energy in a perched state through passive engagement. In birds’ feet, this is facilitated by the digital tendon locking mechanism, which acts like a ratchet, where the tendon interlocks with a tendon sheath to maintain a perched state (Quinn and Baumel, 1990).
A servo motor with a power output of 55 Ncm torque at 7.4 V was chosen for the active actuation. For an optimal centre of gravity location, i.e. centred on the Flapper’s vertical axis, the servo-motor was integrated into the middle of the top bar.
A lever with three arms was designed for the servo-motor, with a pull cable attached to each of the three levers (see figure 2), made from sturdy and lightweight nylon fishing rod:
The right lever connects to one of the outer bars (3) to enable the closing of the gripping mechanism once the servo-motor turns counterclockwise.
@@
The left lever connects to the other outer linkage bar (1) and facilitates the opening of the gripping mechanism once the servo-motor turns clockwise.
Figure 4 shows the different positions of the actuation system and grippers during the stages of the closing procedure.
For the passive part, i.e. allowing sustained perching without energy expenditure, a simple ratchet mechanism had been proposed in our previous work (Broers and Armanini, 2022). However, its rudimentary design could only achieve low precision and, crucially, did not allow for a resetting functionality, thus making an automated take-off from a perched condition impossible.
In the new design, significant changes were introduced to address these limitations, and a new custom spring-loaded ratchet mechanism was developed. The new mechanism consists of a ratchet wheel connected to one of the moving linkage bars and an enclosed spring-loaded pawl that is affixed to the top bar. The ratchet wheel was precision CNC-machined from aluminium, enabling good wear resistance and tight tolerances. Its geometry was improved with miniaturised gear teeth, thus maximising gear teeth count for more precise interlocking. Furthermore, the gear tooth flank design was enhanced for the wheel and pawl, enabling a frictionless tightening of the ratchet. The resulting geometry on the ratchet wheel consisted of 29 evenly spaced teeth over a sector of 37.9° and a root diameter of 64 mm. The tooth profile is an approximate right triangle, with an angle between the perpendicular flank and the hypotenuse of 18.3°. Additionally, a fillet radius of 0.15 mm was chosen for the tooth tip and root for manufacturing purposes.
The curved pawl was likewise CNC-machined from aluminium and included a protruded cable attachment point to enable resetting. It is enclosed in a housing manufactured by precise digital light processing and made out of IGUS Iglidur i3000-PR, a material optimised for sliding parts. Within the housing, a compressed spring constantly pushes the sliding pawl onto the ratchet wheel and therefore allows for the servo-actuated resetting functionality: if the pawl is pulled into the housing, it further compresses the spring and disengages from the ratchet wheel (see figure 5).
@@
Actuating the gripping mechanism’s closure and thus grasping a perching object is facilitated through a custom trigger system: A trigger fork establishes contact with the perching object and activates a mechanical switch housed in the top bar of the linkage. This basic design was kept from the previous prototype (Broers and Armanini, 2022) but improved upon to account for a flapping-wing platform: Its housing features tighter tolerances and a better attachment to the mechanical switch, allowing for more reliable triggering and higher stability in the perched state. Additionally, the required actuation force of the switch was changed from 2.6 N to 1.6 N to accommodate a lighter drone and enable triggering even at lower impact velocities. The selection of the switch model determines the sensitivity of the perching system and, thus, the lower limit of perching contact forces required to initiate the grasping.